# Quantum critical theories in a periodic potential: strange metallic thermoelectric and magneto-transport

### v1

@@ 1 Introduction
Quantum critical theories, where the critical point exists at $T=0$, have the added factor that the critical behavior extends fan-wise into the phase diagram at finite temperature $T$ sachdev_quantum_2023. The temperature $T$ acts as an IR regulator, and probes of the system at a scale less than $T$ are controlled by collective hydrodynamic rather than pure scale- and collisionless behavior. Importantly, the entrance to the hydrodynamic regime halts the RG flow, and whatever remains of the lattice is imprinted in the response. A striking example is a weak chemical potential lattice, $\mu(x)=\bar{\mu}+\delta\mu\cos(Gx)$, with lattice momentum $G<\bar{\mu}$ balm_t-linear_2023: in this case the IR is described by Umklapp hydrodynamics based on Bloch waves, rather than conventional hydrodynamics based on Fourier modes. When instead $G>\bar{\mu}$, the conventional notion that the lattice is irrelevant holds.
@@
Chesler, Lucas, and Sachdev realized that the study of the strong lattice regime $\mu\ll\delta\mu$ in such critical theories becomes tractably simpler by taking the extremal value $\bar{\mu}=0$ chesler_conformal_2014. This eliminates a background scale, simplifying the structure of the solutions and opening the possibility of universal transport regimes exhibiting scaling collapse. This may therefore serve as a testing ground for strongly correlated physics near or at quantum criticality where translational symmetry breaking is strong. In this article we extend their findings of electronic transport in this special regime in 1D lattices to 2D lattices including DC and AC thermal-, thermoelectric- and magneto-transport. As Chesler, Lucas and Sachdev, we use holographic models of quantum criticality to compute these responses.
@@
For an experimental realization of such physics, graphene at charge neutrality naturally comes to mind. Indeed, many studies have considered the scenario of a spatially periodic chemical potential imposed on $\bar{\mu}=0$ charge neutral graphene either in terms of large scale ripples (a 1D periodic potential) or an egg-carton landscape (a 2D periodic potential), see e.g. parkAnisotropicBehaviorsMassless2008; hoSemimetalicGrapheneModulated2009; parkNewGenerationMassless2008; wangElectronicBandGaps2010; bursetTransportSuperlatticesSingle2011; yankowitzEmergenceSuperlatticeDirac2012; breyNonlocalQuantumEffects2020; fernandesEffectiveMediumModel2023. Most of these take a non- or weakly-interacting point of view, even though charged excitations in graphene are arguably not weakly coupled. For 1D potentials, these weak coupling results agree qualitatively with the strong coupling approach in chesler_conformal_2014. We shall show that this is not so for a 2D periodic potential, where the weak coupling approach surprisingly predicts no change to electrical transport bursetTransportSuperlatticesSingle2011. Our approach at strong coupling does however show distinct physics that also differs from the 1D case; thus, both interactions and dimensionality matter.
@@
Due to its quantum critical-like physics at charge neutrality, graphene is in fact an “almost strange metal” sachdevQuantumCriticalityBlack2009; mullerGrapheneNearlyPerfect2009. Our findings are particularly relevant in light of the notion that the phenomenology of the strange metal phase in high-$T_{c}$ cuprate superconductors is also described by such quantum critical physics in the presence of a 2D lattice zaanen_holographic_2015; hartnoll_holographic_2018. This was recently exemplified by computations in a 2D Yukawa-SYK model, which has a holographic dual. This model reproduces the the observed linear-in-$T$-DC conductivity, an AC conductivity with apparent $\omega^{-2/3}$-mid-IR scaling, $T\ln T$ specific heat, Homes’ law relating the $T=0$ superfluid density to $T_{c}\sigma_{DC}(T_{c})$, and broad ARPES width from marginal-Fermi-liquid self-energies $\Sigma(\omega)\sim\omega\ln\omega$ li_strange_2024. The presence of translational symmetry breaking — in li_strange_2024 mainly through disorder in addition to the lattice — is paramount in order to obtain finite and properly scaling transport properties. However, one important phenomenological observation is hitherto unexplained: the magneto-transport. No models as of yet are able to convincingly explain the observed Hall angle scaling $\frac{\sigma_{xx}}{\sigma_{xy}}\sim T^{2}$. In systems dominated by a single momentum relaxation rate, this rate determines both the Hall angle and DC resistivity and these must scale the same way with temperature. Anderson already pointed out that to explain this observed difference in the cuprate strange metal one needs a second relaxation rate, which is not trivial andersonHallEffectTwodimensional1991a. As one of us has recently shown chagnet_natural_2024, by extending Anderson’s earlier arguments to a full hydrodynamic analysis, this requires a system where momentum relaxation through lattice Umklapp or disorder is not the single parametrically small relaxation rate, independently of whether the microscopics is strongly or weakly coupled.
@@
Experiments on certain cuprates (and pnictides) also reveal a longitudinal magnetoresistance on the form hayes_scaling_2016; ayres_incoherent_2021; giraldo-gallo_scale-invariant_2018
@@
where $\alpha$ and $\gamma$ are two dimensionless parameters. This leads to an unsaturating $B$-linear magnetoresistance at large magnetic fields, which stands similarly in stark contrast to the $B^{2}$-scaling of conventional Fermi liquids that saturates to a constant at large fields kittel_introduction_2005. $B$-linear magnetoresistance has been shown to be a general feature of systems with large fluctuations in the conductivity, extending down to moderate magnetic fields in the case of strong inhomogeneities parish_non-saturating_2003. In that case, such systems can be modeled as random resistor networks, which has been shown to be equivalent to “Effective Medium Theory” (EMT) ramakrishnan_equivalence_2017. Within EMT, one considers the macroscopic response of a system constituting of multiple subdomains with varying conductive properties, and may be viewed as an extreme form of translational symmetry breaking. The case of two different types of subdomains at an equal area fraction yields an exactly $B$-linear magnetoresistance guttal_model_2005. Whether all strange metals exhibit a magnetoresistance on quadrature form as in 1 is questioned boyd_single-parameter_2019; nevertheless, a roughly $B$-linear, unsaturating magnetoresistance appears to be a robust feature.
@@
Both of these magnetotransport observations suggest that systems with strong translational symmetry breaking — whether due to an underlying lattice or disorder — should be considered to explain the phenomenology of cuprate strange metals. This provides an additional motivation to study quantum-critical systems in the charge-neutral regime, where the spatially averaged chemical potential vanishes ($\bar{\mu}=0$) but fluctuates locally, $\delta\mu(\vec{x})\neq 0$, beyond purely the fundamental theoretical interest in its universal response. Moreover, at charge neutrality, thermal and electrical transport naturally decouple. In such regimes there are therefore two different independent transport mechanisms that are important in the IR. The resulting universal features are most naturally captured within a holographic framework, which generalizes Landau–Ginzburg theory rather than relying on a specific SYK-like model zaanen_holographic_2015. This is the analysis we pursue in this paper. We provide a complete analysis of the thermoelectric response, both AC and DC, numerically computed, for quantum critical conformal field theories subject to a spatially modulating chemical potential with zero average. In particular, we consider both 1D (Section 4) and 2D (Section 5) “lattices” to find that the dimension of the lattice plays an important role, altering the IR fixed point of the theory, even though both are ultra-local theories evading any RG-like $c$-theorem as found by chesler_conformal_2014. We find that electric transport is completely incoherent, whereas the thermal transport is predominantly coherent and Drude-like. Turning on a magnetic field, we find preliminary evidence of $B$-linear magnetotransport at large fields. Our setup naturally suggests connections to Effective Medium Theory descriptions, and in Section 6 we conclude with an outlook how a combination of this simplified regime with one dimensionless parameter voided provides an obvious direction for further study.
@@ 2 Review: holographic model of strongly correlated quantum criticality
$S=\frac{1}{2\kappa^{2}}\int\,\mathrm{d}^{4}x\,\sqrt{-g}\left[R-2\Lambda-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\right].$ (2)
@@
We employ the de Turck trick to make the equations elliptic and the boundary value problem well-defined by substituting the Ricci tensor with the Harmonic Ricci tensor: $R^{H}_{\mu\nu}=R_{\mu\nu}-\nabla_{(\mu}\xi_{\nu)}$} headrick_new_2010. This imposes the gauge $\xi^{\mu}=0$}, where $\xi^{\mu}=g^{\lambda\nu}(\Gamma^{\mu}_{\lambda\nu}-\tilde{\Gamma}^{\mu}_{\lambda\nu})$} is the de Turck vector and $\tilde{\Gamma}^{\mu}_{\nu\rho}$ is the Christoffel symbol for a suitably chosen background reference metric, which we take to be the analytical solution for an isotropic system ($\delta\mu=0$). In the absence of a background magnetic field, this reference geometry is the AdS4-Schwarzschild solution, whereas for finite magnetic fields we consider the magnetic Reissner-Nordström black brane. We also add a similar gauge-fixing term to Maxwell equations in 4, enforcing $\varphi=0$, where $2\bar{A}=B(x\,\mathrm{d}y-y\,\mathrm{d}x)$} donos_minimally_2016; donos_holographic_2020. This ensures that one can consistently periodically identify the spatial directions $x$ and $y$ as is necessary for the numerics, and imposes a modified Lorenz gauge for the periodic part $A_{\mu}-\bar{A}_{\mu}$ of the background gauge field.
@@
where we have set the black hole horizon location $z_{h}=1$. When $B=0$, we reduce the complexity of the ansatz by directly setting $Q_{ti}=a_{i}=0$}, where $i=\{x,y,z\}$}. [^2] For 1D lattices we also drop the $y$-dependence and set $Q_{xy}=Q_{yz}=0$. In the UV ($z=0$), we demand that the metric asymptotes to AdS and that the time component of the gauge field equals the spatially modulating chemical potential, i.e.,
@@
For nearly all the data presented, we limit ourselves to the universal regime where $\bar{\mu}=0$. The IR boundary conditions ($z=z_{h}$) follow from algebraically solving the equations of motion on the horizon, which leads to the Dirichlet condition $(Q_{tt}-Q_{zz})|_{z=z_{h}}=0$ and Robin-type boundary conditions for the other fields. The resulting non-linear equations of motion are solved numerically on the domain $x,y\in[0,2\pi/G)$, $z\in[0,1]$ with periodic boundary conditions in the transverse $x$ and $y$ directions. Further details on the numerical implementation can be found in Appendix A.
@@ 2.1 Extracting quantum critical responses
Following Refs. donos_thermoelectric_2014; donos_thermoelectric_2015; banks_thermoelectric_2015; donos_holographic_2017, we obtain the DC thermoelectric transport coefficients in the dual quantum critical theory by solving a Stokes flow problem on the horizon on the black hole for a given bulk geometry. This reduces the number of degrees of freedom needed to be solved for as compared to an AC calculation, drops the degree of the PDE by one, and avoids the need to take the numerical $\omega\to 0$} limit of AC data. The Stokes equations read banks_thermoelectric_2015; balm_t-linear_2023
@@
where indices are raised and lowered with respect to the horizon metric $h_{ij}=g_{ij}|_{z=z_{h}}$. The key idea of banks_thermoelectric_2015 is that the spatial averages $\bar{J}^{i}$ and $\bar{Q}^{i}$ do not renormalize as one moves from the horizon to the boundary. The DC thermoelectric response functions in the boundary theory may therefore be extracted from
@@
In our numerics, we fix either the electric field $E^{i}$ or the thermal gradient $\zeta^{i}=-\nabla^{i}T/T$} and study the extracted electric and heat currents $J^{i}$ and $Q^{i}$. In the presence of a nonzero magnetic field $B$, the boundary electric and heat currents decompose into transport and magnetization pieces where the latter should be subtracted in order to obtain the physical transport currents that couple to the external electric field or thermal gradient hartnoll_theory_2007; blake_magnetothermoelectric_2015. However, the magnetization currents vanish at the horizon donos_dc_2016, and as a result, the horizon currents computed via the Stokes flow method directly yield the physical DC conductivities.
@@ 2.2 Extracting AC transport
imposing Lorenz gauge $\nabla^{\nu}\delta A_{\nu}=0$. This ensures that each of the fields in the ansatz below has a kinetic term (i.e., the equations of motion contain a term $\partial_{z}^{2}\delta\Phi$ for all perturbations $\delta\Phi$). We make an ansatz for the perturbed metric and gauge field as
@@
$\sigma^{ij}(\omega)$ $=\frac{1}{i\omega}\Big[G^{R}_{J^{i}J^{j}}(\omega,k=0)-\operatorname{Re}G^{R}_{J^{i}J^{j}}(\omega=0,k\to 0)\Big]\,$, (20)
$T\alpha^{ij}(\omega)$ $=\frac{1}{i\omega}\left[G^{R}_{J^{i}Q^{j}}(\omega,k=0)-\operatorname{Re}G^{R}_{J^{i}Q^{j}}(\omega=0,k\to 0)\right]\,$, (21)
$T\bar{\kappa}^{ij}(\omega)$ $=\frac{1}{i\omega}\left[G^{R}_{Q^{i}Q^{j}}(\omega,k=0)-\operatorname{Re}G^{R}_{Q^{i}Q^{j}}(\omega=0,k\to 0)\right]\,$, (22)
and average the results over the unit cell. The subtraction of the zero-frequency real part removes potential contact terms kovtun_lectures_2012; davison_dissecting_2015. We restrict the AC analysis to zero background magnetic field, for which magnetization currents vanish and no subtraction is required.

### v2

# Quantum critical theories in a periodic potential: strange metallic thermoelectric and magnetotransport

@@ 1 Introduction
Quantum critical theories, where the critical point exists at $T=0$, have the added factor that the critical behavior extends fan-wise into the phase diagram at finite temperature $T$ Sachdev (2023). The temperature $T$ acts as an IR regulator, and probes of the system at an energy scale less than $T$ are controlled by collective hydrodynamic rather than pure scale-invariant and collisionless behavior. Importantly, the entrance to the hydrodynamic regime halts the RG flow, and whatever remains of the lattice is imprinted in the response. A striking example is a weak chemical potential lattice, $\mu(x)=\bar{\mu}+\delta\mu\cos(Gx)$, with lattice momentum $G<\bar{\mu}$ Balm et al. (2023): in this case the IR is described by Umklapp hydrodynamics based on Bloch waves, rather than conventional hydrodynamics based on Fourier modes. When instead $G>\bar{\mu}$, the conventional notion that the lattice is irrelevant holds.
@@
Chesler, Lucas, and Sachdev realized that the study of the strong lattice regime $\delta\mu\gg\bar{\mu}$ in such critical theories becomes tractably simpler by taking the extremal value $\bar{\mu}=0$ Chesler et al. (2014). This eliminates a background scale, simplifying the structure of the solutions and opening the possibility of universal transport regimes exhibiting scaling collapse. This may therefore serve as a testing ground for strongly correlated physics near or at quantum criticality where translational symmetry breaking is strong. In this article we extend their findings of electronic transport in this special regime from 1D to 2D lattices, covering DC and AC thermal, thermoelectric, and magnetotransport. As Chesler, Lucas and Sachdev, we use holographic models of quantum criticality to compute these responses.
@@
For an experimental realization of such physics, graphene at charge neutrality naturally comes to mind. Indeed, many studies have considered the scenario of a spatially periodic chemical potential imposed on $\bar{\mu}=0$ charge neutral graphene either in terms of large scale ripples (a 1D periodic potential) or an egg-carton landscape (a 2D periodic potential), see e.g., Park et al. (2008a); Ho et al. (2009); Park et al. (2008b); Wang and Zhu (2010); Burset et al. (2011); Yankowitz et al. (2012); Brey et al. (2020); Fernandes (2023). Most of these adopt a non-interacting or weakly interacting point of view, even though charged excitations in graphene are arguably not weakly coupled. For 1D potentials, these weak coupling results agree qualitatively with the strong coupling approach in Chesler et al. (2014). We shall show that this is not so for a 2D periodic potential, where the weak coupling approach surprisingly predicts no change to electrical transport Burset et al. (2011). Our approach at strong coupling does however show distinct physics that also differs from the 1D case; thus, both interactions and dimensionality matter.
@@
Due to its quantum critical-like physics at charge neutrality, graphene is in fact an “almost strange metal” Sachdev and Mueller (2009); Müller et al. (2009). Our findings are particularly relevant for the truly strange metal phase in high-$T_{c}$ cuprate superconductors in light of the hypothesis that its phenomenology is described by such quantum critical physics in the presence of a 2D lattice; see e.g., Zaanen et al. (2015); Hartnoll et al. (2018) for a review. This was recently exemplified by computations in a 2D Yukawa-SYK model, which has a holographic dual. This model reproduces the the observed linear-in-$T$ DC conductivity, an AC conductivity with apparent $\omega^{-2/3}$ mid-IR scaling, $T\ln T$ specific heat, Homes’ law relating the $T=0$ superfluid density to $T_{c}\sigma_{DC}(T_{c})$, and broad ARPES width from marginal Fermi liquid self-energies $\Sigma(\omega)\sim\omega\ln\omega$ Li et al. (2024). The presence of translational symmetry breaking — in Li et al. (2024) mainly through disorder in addition to the lattice — is paramount in order to obtain finite and properly scaling transport properties. However, one important phenomenological observation is hitherto unexplained: the magnetotransport. No models as of yet are able to convincingly explain the observed Hall angle scaling $\frac{\sigma_{xx}}{\sigma_{xy}}\sim T^{2}$}. In systems dominated by a single momentum relaxation rate, this rate determines both the Hall angle and DC resistivity and these must scale the same way with temperature. Anderson already pointed out that to explain this observed difference in the cuprate strange metal one needs a second relaxation rate, which is not trivial Anderson (1991). As one of us has recently shown Chagnet et al. , extending Anderson’s earlier arguments to a full hydrodynamic analysis requires a system in which momentum relaxation through lattice Umklapp or disorder is not the only parametrically small relaxation rate, independently of whether the microscopics is strongly or weakly coupled. An important ingredient in reaching this conclusion is the AC response, in addition to the DC response. While analyses based solely on DC transport have been remarkably successful in describing strange metals Amoretti et al. (2020); Blake and Donos (2015), such theories typically rely on a separation of mechanisms: the longitudinal DC resistivity is assumed to be of quantum critical origin, whereas the Hall conductivity is not. This effectively corresponds to scenarios in which the system behaves as if it were charge neutral, or exhibits strong translational symmetry breaking for longitudinal transport but weak symmetry breaking at finite charge density for Hall transport. This picture is prima facie in tension with measurements of the AC longitudinal optical conductivity, which display a clear Drude peak at low temperatures. The search for systems that exhibits both Drude-like AC optical conductivity and anomalous Hall transport therefore motivates our study of systems in the extreme limit of strong translational symmetry breaking.
@@
Experiments on certain cuprates (and pnictides) also reveal a longitudinal magnetoresistance on the form Hayes et al. (2016); Ayres et al. (2021); Giraldo-Gallo et al. (2018)
@@
where $\alpha$ and $\gamma$ are two dimensionless parameters. This leads to an unsaturating $B$-linear magnetoresistance at large magnetic fields, which stands similarly in stark contrast to the $B^{2}$-scaling of conventional Fermi liquids that saturates to a constant at large fields Kittel (2005). $B$-linear magnetoresistance has been shown to be a general feature of systems with large spatial fluctuations in the conductivity, extending down to moderate magnetic fields in the case of strong inhomogeneities Parish and Littlewood (2003). In that case, such systems can be modeled as random resistor networks, which has been shown to be equivalent to “Effective Medium Theory” (EMT) Ramakrishnan et al. (2017). Within EMT, one considers the macroscopic response of a system constituting of multiple subdomains with varying conductive properties, and may be viewed as an extreme form of translational symmetry breaking. The case of two different types of subdomains at an equal area fraction yields an exactly $B$-linear magnetoresistance Guttal and Stroud (2005). [^1] In holographic approaches there exist so-called DBI models which predict a resistivity of the form 1 Kiritsis and Li (2017), but the physical origin is an underlying relativistic symmetry rather than an averaging. Whether all strange metals exhibit a magnetoresistance on quadrature form as in 1 is questioned Boyd and Phillips (2019); nevertheless, a roughly $B$-linear, unsaturating magnetoresistance appears to be a robust feature.
@@
Both of these magnetotransport observations suggest that systems with strong translational symmetry breaking — whether due to an underlying lattice or disorder — should be considered to explain the phenomenology of cuprate strange metals. This provides an additional motivation to study quantum-critical systems in the charge-neutral regime, where the spatially averaged chemical potential vanishes ($\bar{\mu}=0$) but fluctuates locally, $\delta\mu(\vec{x})\neq 0$, beyond purely the fundamental theoretical interest in its universal response. Moreover, at charge neutrality, thermal and electrical transport naturally decouple. In such regimes there are therefore two different independent transport mechanisms that are important in the IR. The resulting universal features are most naturally captured within a holographic framework, which generalizes Landau–Ginzburg theory rather than relying on a specific SYK-like model Zaanen et al. (2015). This is the analysis we pursue in this paper. We provide a complete analysis of the thermoelectric response, both AC and DC, numerically computed, for quantum critical conformal field theories subject to a spatially modulating chemical potential with zero average. In particular, we consider both 1D (Section 4) and 2D (Section 5) “lattices” to find that the dimension of the lattice plays an important role, altering the IR fixed point of the theory, even though both are ultra-local theories evading any RG-like $c$-theorem as found by Chesler et al. (2014). We find that electric transport is completely incoherent, whereas the thermal transport is predominantly coherent and Drude-like. Turning on a magnetic field, we find preliminary evidence of $B$-linear magnetotransport at large fields. Our setup naturally suggests connections to Effective Medium Theory descriptions, and in Section 6 we conclude with an outlook how a combination of this simplified regime with one dimensionless parameter voided provides an obvious direction for further study.
@@ 2 Review: holographic model of strongly correlated quantum criticality
$S=\frac{1}{2\kappa^{2}}\int\,\mathrm{d}^{4}x\,\sqrt{-g}\left[R-2\Lambda-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\right]\,$, (2)
@@
We employ the de Turck trick to make the equations elliptic and the boundary value problem well-defined by substituting the Ricci tensor with the Harmonic Ricci tensor: $R^{H}_{\mu\nu}=R_{\mu\nu}-\nabla_{(\mu}\xi_{\nu)}$} Headrick et al. (2010). This imposes the gauge $\xi^{\mu}=0$}, where $\xi^{\mu}=g^{\lambda\nu}(\Gamma^{\mu}_{\lambda\nu}-\tilde{\Gamma}^{\mu}_{\lambda\nu})$} is the de Turck vector and $\tilde{\Gamma}^{\mu}_{\nu\rho}$ is the Christoffel symbol for a suitably chosen background reference metric, which we take to be the analytical solution for an isotropic system ($\delta\mu=0$). In the absence of a background magnetic field, this reference geometry is the AdS4-Reissner-Nordström solution, whereas for finite magnetic fields we consider the dyonic Reissner-Nordström black brane. We also add a similar gauge-fixing term to Maxwell’s equations in 4, enforcing $\varphi=0$, where $2\bar{A}=B(x\,\mathrm{d}y-y\,\mathrm{d}x)$} Donos and Gauntlett (2016); Donos et al. (2020). This ensures that one can consistently periodically identify the spatial directions $x$ and $y$ as is necessary for the numerics, and imposes a modified Lorenz gauge for the periodic part $A_{\mu}-\bar{A}_{\mu}$ of the background gauge field.
@@
where we have set the black hole horizon location $z_{h}=1$. For nearly all the data presented, we limit ourselves to the universal regime where $\bar{\mu}=0$. When $B=0$, we reduce the complexity of the ansatz by directly setting $Q_{ti}=a_{i}=0$}, where $i=\{x,y,z\}$}. [^3] For 1D lattices we also drop the $y$-dependence and set $Q_{xy}=Q_{yz}=0$. In the UV ($z=0$), we demand that the metric asymptotes to AdS4 and that the time component of the gauge field equals the spatially modulating chemical potential, i.e.,
@@
The IR boundary conditions ($z=z_{h}$) follow from algebraically solving the equations of motion on the horizon, which leads to the Dirichlet condition $(Q_{tt}-Q_{zz})|_{z=z_{h}}=0$ and Robin-type boundary conditions for the other fields. The resulting non-linear equations of motion are solved numerically on the domain $x,y\in[0,2\pi/G)$, $z\in[0,1]$ with periodic boundary conditions in the transverse $x$ and $y$ directions. Further details on the numerical implementation can be found in Appendix A.
@@ 2.1 Extracting quantum critical responses
Following Refs. Donos and Gauntlett (2014b, 2015); Banks et al. (2015); Donos et al. (2017), we obtain the DC thermoelectric transport coefficients in the dual quantum critical theory by solving a Stokes flow problem on the horizon on the black hole for a given bulk geometry. This reduces the number of degrees of freedom needed to be solved for as compared to an AC calculation, drops the degree of the PDE by one, and avoids the need to take the numerical $\omega\to 0$} limit of AC data. The Stokes equations read Banks et al. (2015); Balm et al. (2023)
@@
where indices are raised and lowered with respect to the horizon metric $h_{ij}=g_{ij}|_{z=z_{h}}$. The key idea of Banks et al. (2015) is that the spatial averages $\bar{J}^{i}$ and $\bar{Q}^{i}$ (taken over the unit cell) do not renormalize as one moves from the horizon to the boundary. The DC thermoelectric response functions in the boundary theory may therefore be extracted from
@@
In our numerics, we fix either the electric field $E^{i}$ or the thermal gradient $\zeta^{i}=-\nabla^{i}T/T$} and study the extracted electric and heat currents $J^{i}$ and $Q^{i}$. In the presence of a nonzero magnetic field $B$, the boundary electric and heat currents decompose into transport and magnetization pieces where the latter should be subtracted in order to obtain the physical transport currents that couple to the external electric field or thermal gradient Hartnoll et al. (2007); Blake et al. (2015). However, the magnetization currents vanish at the horizon Donos et al. (2016), and as a result, the horizon currents computed via the Stokes flow method directly yield the physical DC conductivities.
@@ 2.2 Extracting AC transport
imposing the Lorenz gauge $\nabla^{\nu}\delta(A_{\nu})=0$. This ensures that each of the fields in the ansatz below has a kinetic term (i.e., the equations of motion contain a term $\partial_{z}^{2}\delta\Phi$ for all perturbations $\delta\Phi$). We make an ansatz for the perturbed metric and gauge field as
@@
$\sigma^{ij}(\omega)$ $=\frac{i}{\omega}\Big[G^{R}_{J^{i}J^{j}}(\omega,k=0)-\operatorname{Re}G^{R}_{J^{i}J^{j}}(\omega=0,k\to 0)\Big]\,$, (20)
$T\alpha^{ij}(\omega)$ $=\frac{i}{\omega}\left[G^{R}_{J^{i}Q^{j}}(\omega,k=0)-\operatorname{Re}G^{R}_{J^{i}Q^{j}}(\omega=0,k\to 0)\right]\,$, (21)
$T\bar{\kappa}^{ij}(\omega)$ $=\frac{i}{\omega}\left[G^{R}_{Q^{i}Q^{j}}(\omega,k=0)-\operatorname{Re}G^{R}_{Q^{i}Q^{j}}(\omega=0,k\to 0)\right]\,$, (22)
and average the results over the unit cell. The subtraction of the zero-frequency real part removes potential contact terms Kovtun (2012); Davison and Goutéraux (2015). We restrict the AC analysis to zero background magnetic field, for which magnetization currents vanish and no subtraction is required.